Can someone help me solve the exponential equation, please?
9^8x = 27

Answer by Alan3354(69443)

9^8x = 27
8x*log(9) = log(27)
x = log(27)/(8log(9)
-------------
Or
3^16x = 3^3
16x = 3
etc